How much ice at 0°C must be added to 100 g water at 30°C in order to reduce its temperature to 20°C?
(The latent heat of fusion of water is 80 cal/g)

Grade:12th Pass

2 Answers

melvin davis
13 years ago

heat lost by the water = heat gained by ice for changing into water and rising the temp. to 20degree celsius

100*1*(30-20) = m*80 + m*1*(20-0)

solve it

Sudheesh Singanamalla
13 years ago

specific heat capacity of water = 1 cal / g ; Tf = 20 ; Ti = 30 ;

Heat lost by water = m * s * (Tf - Ti )

= 100 * 1 cal/g * ( 20 - 30 ) = -1000 J

Heat required to melt ice = m L = m * 80

Heat required to raise temperature of ice water to 20 = m * 1 cal/g * 10

heat lost = heat gained

-1000 = -m * 80 + 10 * m

-1000 = -90 * m

mass of ice (m) = 1000/90 = 100/9 = 11.11 grams.

 

Therefore 1.11 grams of ice is required to change water from 30 degrees to 20 degree celcius.

Please Approve !